Calculate the percent yield for the reaction: P4 (s) + 6 Cl2 (g) → 4 PCl3 (l) if 75.0 g of phosphorus reacts with excess chlorine gas to produce 111.0 g of phosphorus trichloride.

1
Expert's answer
2022-04-16T04:09:28-0400

P4 (s) + 6 Cl2 (g) → 4 PCl3 (l)

According to the reaction, n (PCl3) = 4 x n (P4).

n = m / M

M (P4) = 123.9 g/mol

M (PCl3) = 137.3 g/mol

n (P4) = 75/123.9 = 0.61 mol

n (PCl3) = 4 x n (P4) = 4 x 0.61 = 2.4 mol

m (PCl3)theoretical = n x M = 2.4 x 137.3 = 332.4 g

%(PCl3) = 111 / 332.4 x 100 = 33.4%